# Algorithms for Approximating Conditionally Optimal Bounds

### v2

@@ 1 Preliminaries
Suppose for the sake of contradiction that ${\bm{x}}<_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$. Since ${\bm{x}}<_{R}{\bm{y}}$ we know that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$. But since ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$ we have by Definition 1 that ${\bm{x}}\sim_{R}{\bm{y}}$, which contradicts the earlier observation that ${\bm{x}}\not\sim_{R}{\bm{y}}$. ∎
@@ 3 Conditional-Optimality
Total order $T$ agrees with a preorder $R$ if $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}<_{R}{\bm{y}}\Rightarrow{\bm{x}}<_{T}{\bm{y}}$.
@@
A bound $B$ is consistent with preorder $R$ if it is consistent with every total order that agrees with $R$.
@@
Suppose $B$ is consistent with preorder $R$ and take first any ${\bm{x}},{\bm{y}}\in\Omega$ where ${\bm{x}}<_{R}{\bm{y}}$. Definition 13 dictates that $B$ must be consistent with every total order $T$ agreeing with $R$. And Definition 12 requires that we must also have ${\bm{x}}<_{T}{\bm{y}}$, which in turn implies that $B({\bm{x}})\leq B({\bm{y}})$ by the definition of consistency for total orders. Next, take ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$. Let $T$ be any total order agreeing with $R$ such that ${\bm{x}}<_{T}{\bm{y}}$, and let $T^{\prime}$ be the total order identical to $T$ except that it reverses the order of $\bm{x}$} and $\bm{y}$}. Since $B$ is consistent with $R$, Definition 13 ensures that $B$ is consistent with both $T$ and $T^{\prime}$. This implies that $B({\bm{x}})\leq B({\bm{y}})$ and $B({\bm{y}})\leq B({\bm{x}})$ so that $B({\bm{x}})=B({\bm{y}})$.
Now suppose $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$ and $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ for preorder $R$. Let $T$ be any total order agreeing with $R$. For each ${\bm{x}},{\bm{y}}\in\Omega$ we have by Proposition 1 and taking the contrapositive of Definition 12 that
@@
If ${\bm{x}}<_{R}{\bm{y}}$, then by assumption $B({\bm{x}})\leq B({\bm{y}})$. And if ${\bm{x}}\sim_{R}{\bm{y}}$, we again have by assumption that $B({\bm{x}})=B({\bm{y}})\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Thus, in any case, ${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$, which implies, by the definition of consistency for total orders, that $B$ is consistent with $T$. Since $T$ was an arbitrary total order agreeing with $R$, we have by Definition 13 that $B$ is also consistent with $R$. ∎
@@
We begin by noting that, according to Lemma 4, $B_{R}^{*}$ is consistent with $R$. Thus, by Definition 14, it suffices to show that for any valid bound $B_{R}$, consistent with $R$, and $\forall{\bm{x}}\in\Omega,B^{*}_{R}({\bm{x}})\geq B_{R}({\bm{x}})$. This would imply that $B_{R}$ is either strictly worse than $B_{R}^{*}$ for some ${\bm{x}}\in\Omega$, or that $B_{R}=B_{R}^{*}$. To that end, suppose to the contrary that there exists some valid $B_{R}$, also consistent with $R$, such that $\exists{\bm{x}}\in\Omega:B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. Now let $T$ be any total order agreeing with $R$ such that $\Omega({\bm{x}},T)=\Omega({\bm{x}},R)$. Since $\Omega({\bm{x}},T)=\Omega({\bm{x}},R)$ we also have by Definition 15 that $B_{T}^{*}({\bm{x}})=B_{R}^{*}({\bm{x}})$. So it must be the case that $B_{T}^{*}({\bm{x}})<B_{R}({\bm{x}})$ as well, which is a contradiction given that $B_{T}^{*}$ is conditionally optimal by Theorem 1. ∎
@@ 5.1 Calculating bounds for homogeneous samples
From Learned-Miller [2] Lemma 2.2, the result is true for $S_{i}=S_{\min}$, so we fix $\bm{S}_{i}\neq\bm{S}_{\min}$. By Theorem 1 we need only show that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}=\mu$ falls within the interval defined by Inequality 28. To that end, by Proposition 6, it will suffice to show the following two properties. On one hand, $\forall H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}$, $E[H]>S_{\min}\sqrt[n]{1-\alpha}+S_{i}(1-\sqrt[n]{1-\alpha})$. And on the other hand, there exists a sequence of distributions in $\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ whose limiting mean is bounded above by $S_{\min}\sqrt[n]{1-\alpha}+S_{i+1}(1-\sqrt[n]{1-\alpha})$.
@@ 6 Refinements of $\mathcal{F}$
Also by construction and Definition 9, $P_{G}[x_{(i)}\leq X_{(i)}]=P_{H_{p}}[x_{(i)}\leq X_{(i)}]$. Therefore, $H_{p}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$. Take $H_{p^{*}}$ to be the distribution in $\mathcal{F}_{C^{-}}$ corresponding to $G^{*}$. By letting $\varepsilon\rightarrow 0$ and by Inequality 35 and the definition of $G^{*}$ we have
@@
From Lemma 9 it will suffice to show that for any two distributions $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},T_{\ell})]=P_{H}[\Omega({\bm{x}},T_{\ell})]$. According to Definition 7, we have that $G\in\mathcal{G}(\Omega(\bm{x},T_{\ell}),\alpha)\}\Leftrightarrow P_{G}[L_{1}({\bm{x}},{\bm{Y}})]>\alpha$. But, for each $j\in[n]$ and any ${\bm{x}},{\bm{y}}\in\Omega$, the events $(x_{(j)}<y_{(j)})$ and $(x_{(j)}=y_{(j)})$ are mutually exclusive, so from Equation 5 we have that for all $j\in[n]$

### v3

@@ 1 Preliminaries
To prove the forward direction, suppose for the sake of contradiction that ${\bm{x}}<_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$. Since ${\bm{x}}<_{R}{\bm{y}}$ we know that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$. But since ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$ we have by Definition 1 that ${\bm{x}}\sim_{R}{\bm{y}}$, which contradicts the earlier observation that ${\bm{x}}\not\sim_{R}{\bm{y}}$.
We prove the contrapositive of the reverse direction: ${\bm{y}}\lesssim_{R}{\bm{x}}\Rightarrow\neg({\bm{x}}<_{R}{\bm{y}})$. Suppose again for the sake of contradiction that ${\bm{y}}\lesssim_{R}{\bm{x}}$ and ${\bm{x}}<_{R}{\bm{y}}$. The latter implies that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$. But together ${\bm{y}}\lesssim_{R}{\bm{x}}$ and ${\bm{x}}\lesssim_{R}{\bm{y}}$ imply by Definition 1 that ${\bm{x}}\not\sim_{R}{\bm{y}}$, a contradiction. ∎
@@ 3 Conditional-Optimality
Preorder order $D$ agrees with a preorder $R$ if $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}<_{R}{\bm{y}}\Rightarrow{\bm{x}}<_{D}{\bm{y}}$. Note that agreement is not necessarily a symmetric property.
Proposition 5.
If total order $T$ agrees with preorder $R$, then $\forall{\bm{x}}\in\Omega$, $\Omega({\bm{x}},T)\subseteq\Omega({\bm{x}},R)$.
Proof.
Let ${\bm{x}}\in\Omega$ be arbitrary and note that by Definition 12, if total order $T$ agrees with preorder $R$, then for each ${\bm{y}}\in\Omega$
$\neg({\bm{y}}<_{T}{\bm{x}})\Rightarrow\neg({\bm{y}}<_{R}{\bm{x}}).$ (13)
It follows that
[…]
Proposition 6.
For every preorder $R$, there exists a total order $T$ that agrees with $R$.
Proof.
We proceed in two stages by first showing that there exists a partial order that agrees with $R$ and then invoking the Szpilrajn extension theorem [3] to argue that there exists a total order that agrees with the partial order.
Every preorder $R$ is a partial order over equivalence classes. Construct partial order $D$ by maintaining all non-equivalence relations of $R$ and arbitrarily assigning a total order between the elements within each of the equivalence classes of $R$. Now take any ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}<_{R}{\bm{y}}$. Since ${\bm{x}}\not\sim_{R}{\bm{y}}$, it’s clear by construction that ${\bm{x}}<_{D}{\bm{y}}$ as well. Thus, partial order $D$ agrees with $R$.
The Szpilrajn extension theorem [3] establishes that there must also exist a total order $T$ such that $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\leq_{D}{\bm{y}}\Rightarrow{\bm{x}}\leq_{T}{\bm{y}}$. Fix ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}<_{D}{\bm{y}}$ and let $T$ be the total order guaranteed by the theorem. We must then have that ${\bm{x}}\leq_{T}{\bm{y}}$. And of course ${\bm{x}}\neq{\bm{y}}$ by Definition 2. Thus, by Definition 3, ${\bm{x}}<_{T}{\bm{y}}$ as well. Therefore, total order $T$ agrees with $D$. ∎
Proposition 7.
For every total order $T$ agreeing with preorder $R$ and any ${\bm{x}}\in\Omega$, there exists a ${\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$ and $\Omega({\bm{x}},R)=\Omega({\bm{y}},T)$.
Proof.
Fix ${\bm{x}}\in\Omega$ and take $\bm{y}$} such that ${\bm{x}}\sim_{R}{\bm{y}}$ and $\forall{\bm{z}}\in\Omega:{\bm{x}}\sim_{R}{\bm{z}}$, ${\bm{y}}\leq_{T}{\bm{z}}$, i.e. $\bm{y}$} is the smallest element according to $T$ that is equivalent to $\bm{x}$} according to $R$. Such a $\bm{y}$} must exist by the well-ordering principle. By Definition 4, $\Omega({\bm{x}},R)=\Omega({\bm{y}},R)$ so that, invoking Proposition 5, it suffices to show that $\Omega({\bm{y}},R)\subseteq\Omega({\bm{y}},T)$.
We can partition $\Omega({\bm{y}},R)$ such that $\Omega({\bm{y}},R)=S_{1}\cup S_{2}$ where $S_{1}=\{{\bm{z}}:{\bm{y}}\sim_{R}{\bm{z}}\}$ and $S_{2}=\{{\bm{z}}:{\bm{y}}<_{R}{\bm{z}}\}$. Now take any ${\bm{z}}\in\Omega({\bm{y}},R)$. If ${\bm{z}}={\bm{y}}$, then it’s clear by Definition 4 that ${\bm{z}}\in\Omega({\bm{y}},T)$. If ${\bm{z}}\in S_{1}\setminus\{{\bm{y}}\}$, then ${\bm{y}}<_{T}{\bm{z}}$ by our choice of $\bm{y}$}, so that, again by Definition 4, ${\bm{z}}\in\Omega({\bm{y}},T)$. Finally, suppose that ${\bm{z}}\in S_{2}$. Since $T$ agress with $R$, we have by Definition 12 that ${\bm{y}}<_{T}{\bm{z}}$. Thus, by Definition 4, ${\bm{z}}\in\Omega({\bm{y}},T)$, and therefore $\Omega({\bm{y}},R)\subseteq\Omega({\bm{y}},T)$. ∎
@@
A bound $B$ is consistent with total order $T$ if $\forall{\bm{x}},{\bm{y}}\in\Omega$,
${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}}).$ (15)
Definition 14.
Bound $B$ is consistent with preorder $R$ if it is consistent with every total order that agrees with $R$.
@@
Suppose $B$ is consistent with preorder $R$ and take first any ${\bm{x}},{\bm{y}}\in\Omega$ where ${\bm{x}}<_{R}{\bm{y}}$. Definition 14 dictates that $B$ must be consistent with every total order $T$ agreeing with $R$. And Definition 12 requires that we must also have ${\bm{x}}<_{T}{\bm{y}}$, which in turn implies, by Definition 13, that $B({\bm{x}})\leq B({\bm{y}})$. Next, take ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$. Let $T$ be any total order agreeing with $R$ such that ${\bm{x}}<_{T}{\bm{y}}$, and let $T^{\prime}$ be the total order identical to $T$ except that it reverses the order of $\bm{x}$} and $\bm{y}$}. Since $B$ is consistent with $R$, Definition 14 ensures that $B$ is consistent with both $T$ and $T^{\prime}$. This implies that $B({\bm{x}})\leq B({\bm{y}})$ and $B({\bm{y}})\leq B({\bm{x}})$ so that $B({\bm{x}})=B({\bm{y}})$.
Now suppose that $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$ and $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ for preorder $R$. Let $T$ be any total order agreeing with $R$. For each ${\bm{x}},{\bm{y}}\in\Omega$ we have by Proposition 1 and taking the contrapositive of Definition 12 that
@@
If ${\bm{x}}<_{R}{\bm{y}}$, then by assumption $B({\bm{x}})\leq B({\bm{y}})$. And if ${\bm{x}}\sim_{R}{\bm{y}}$, we again have by assumption that $B({\bm{x}})=B({\bm{y}})\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Thus, in any case, ${\bm{x}}\lesssim_{R}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. It follows then by Implication 16 that ${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Therefore, by Definition 13, it must be the case that $B$ is consistent with $T$. Since $T$ was an arbitrary total order agreeing with $R$, we have, by Definition 14, that $B$ is also consistent with $R$. ∎
@@
We begin by noting that, according to Lemma 4, $B_{R}^{*}$ is consistent with $R$. Thus, by Definition 15, it suffices to show that for any valid bound $B_{R}$, consistent with $R$, and $\forall{\bm{x}}\in\Omega,B^{*}_{R}({\bm{x}})\geq B_{R}({\bm{x}})$. This would imply that $B_{R}$ is either strictly worse than $B_{R}^{*}$ for some ${\bm{x}}\in\Omega$, or that $B_{R}=B_{R}^{*}$. To that end, suppose to the contrary that there exists some valid $B_{R}$, also consistent with $R$, such that $\exists{\bm{x}}\in\Omega:B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. Now let $T$ be any total order agreeing with $R$ and take ${\bm{y}}\in\Omega$ such that ${\bm{y}}\sim_{R}{\bm{x}}$ and $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)=\Omega({\bm{x}},R)$. We are guaranteed that such a $T$ exists by Propositions 6, and we are guaranteed that such a $\bm{y}$} exists by Proposition 7. Since ${\bm{x}}\sim_{R}{\bm{y}}$ and $B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$, it’s clear by Lemma 3 that
$B_{R}^{*}({\bm{y}})=B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})=B_{R}({\bm{y}}).$ (18)
And because $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)$ we also have by Definition 16 that $B_{T}^{*}({\bm{y}})=B_{R}^{*}({\bm{y}})$. So it must be the case that $B_{T}^{*}({\bm{x}})<B_{R}({\bm{y}})$ as well, which is a contradiction given that $B_{T}^{*}$ is conditionally optimal by Theorem 1. ∎
@@ 5.1 Calculating bounds for homogeneous samples
From Learned-Miller [2] Lemma 2.2, the result is true for $S_{i}=S_{\min}$, so we fix $\bm{S}_{i}\neq\bm{S}_{\min}$. By Theorem 1 we need only show that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}=\mu$ falls within the interval defined by Inequality 32. To that end, by Proposition 9, it will suffice to show the following two properties. On one hand, $\forall H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}$, $E[H]>S_{\min}\sqrt[n]{1-\alpha}+S_{i}(1-\sqrt[n]{1-\alpha})$. And on the other hand, there exists a sequence of distributions in $\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ whose limiting mean is bounded above by $S_{\min}\sqrt[n]{1-\alpha}+S_{i+1}(1-\sqrt[n]{1-\alpha})$.
@@ 6 Refinements of $\mathcal{F}$
Also by construction and Definition 9, $P_{G}[x_{(i)}\leq X_{(i)}]=P_{H_{p}}[x_{(i)}\leq X_{(i)}]$. Therefore, $H_{p}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$. Take $H_{p^{*}}$ to be the distribution in $\mathcal{F}_{C^{-}}$ corresponding to $G^{*}$. By letting $\varepsilon\rightarrow 0$ and by Inequality 39 and the definition of $G^{*}$ we have
@@
From Lemma 9 it will suffice to show that for any two distributions $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},T_{\ell})]=P_{H}[\Omega({\bm{x}},T_{\ell})]$. According to Definition 7, we have that $G\in\mathcal{G}(\Omega(\bm{x},T_{\ell}),\alpha)\Leftrightarrow P_{G}[L_{1}({\bm{x}},{\bm{Y}})]>\alpha$. But, for each $j\in[n]$ and any ${\bm{x}},{\bm{y}}\in\Omega$, the events $(x_{(j)}<y_{(j)})$ and $(x_{(j)}=y_{(j)})$ are mutually exclusive, so from Equation 5 we have that for all $j\in[n]$